# Online Bayesian Experimental Design for Partially Observable Dynamical Systems

### v1

@@ 3.3 Monte Carlo estimators using NPF
Thus, to approximate the EIG in (4) we draw: (i) parameters $\{\boldsymbol{\theta}_{t}^{(m)},\dot{\boldsymbol{\theta}}_{t}^{(m)}\}$ for $m=1,\dots,M$; (ii) conditional state trajectories $\{\boldsymbol{x}_{0:t-1}^{(m,n)},\widetilde{\boldsymbol{x}}_{t}^{(m,n)},\dot{\boldsymbol{x}}_{t}^{(m,n)},\ddot{\boldsymbol{x}}_{t}^{(m,n)}\}$ for $n=1,\dots,N$; and (iii) pseudo-observations $\widetilde{\boldsymbol{y}}_{t}^{(\ell)}\!\sim\!g(\boldsymbol{y}_{t}\!\mid\!\widetilde{\boldsymbol{x}}_{t}^{(m,n)},\boldsymbol{\theta}^{(m)},\boldsymbol{\xi}_{t})$, for $\ell=1,\dots,L$. This yields the nested Monte Carlo approximation
@@
Following a similar procedure, the gradients $\nabla_{\boldsymbol{\xi}_{t}}L_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})$ and $\nabla_{\boldsymbol{\xi}_{t}}Z_{\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})$ of (3.2) can be approximated. See details in Appendix LABEL:supp-ap:approximations.
@@ 3.4 Consistency of the EIG estimator
Consistency of our estimator $\widehat{\mathcal{I}}(\boldsymbol{\xi}_{t})$ follows from two ingredients: (i) NPF convergence under standard conditions (small/regular jittering kernels, Lipschitz state posterior in $\boldsymbol{\theta}$, and a bounded positive likelihood), and (ii) NMC consistency provided the EIG integrand, $\log(L/Z)$, is Lipschitz and square-integrable. Formal statements of these assumptions are given in Appendix LABEL:supp-ap:consistency.
Theorem 1.
Let $\widehat{\mathcal{I}}(\boldsymbol{\xi}_{t})$ denote the nested Monte Carlo estimator of the EIG defined in (8), constructed using $M$ parameter particles, $N$ state particles per parameter, and $L$ pseudo-observations. Under Assumptions 1–5 (in Appendix LABEL:supp-ap:consistency), it holds that
$\widehat{\mathcal{I}}(\boldsymbol{\xi}_{t})\;\xrightarrow[L,M,N\to\infty]{\text{a.s.}}\;\mathcal{I}(\boldsymbol{\xi}_{t})$
for any $t$ and $\boldsymbol{\xi}_{t}$.
The proof combines NMC results (Rainforth18), ensuring consistency of nested expectations under Assumption 5, with NPF convergence guarantees (Crisan18bernoulli), which establish almost-sure convergence of empirical state–parameter measures under Assumptions 1–4. Together, these results imply that the Monte Carlo averages in our estimator converge almost surely to the true EIG. Full details are given in Appendix LABEL:supp-ap:consistency.
@@ 3.5 Overall algorithm
Algorithm 1 summarizes our proposed framework, BAD-PODS (Bayesian Adaptive Design for Partially Observable Dynamical Systems). At each time step $t$, the design $\boldsymbol{\xi}_{t}$ is optimized by SGA (Section 3.2), using the gradient estimator of the EIG computed with NMC samples from $\Gamma_{\boldsymbol{\xi}_{t}}$ (Section 3.3). Iterative updates over $K$ steps, with step sizes $\eta_{k}$, yield the optimized design $\boldsymbol{\xi}_{t}^{\star}$, after which the real observation $\boldsymbol{y}_{t}$ is collected. The posterior is then updated via a NPF (Section 2.4): parameter particles are jittered, states are propagated through $f$, and weights are adjusted using $g$, with resampling to maintain diversity. This recursive construction avoids reprocessing past data, with a per-step cost of $\mathcal{O}((KL+1)MN)$ (and $\mathcal{O}(KLMNt)$ over $t$ steps), ensuring linear scaling in time while supporting gradient-based design optimization.
The algorithm takes as input the prior distributions, particle numbers $(M,N)$, and optimization hyperparameters $(K,\eta_{k})$, and produces as output the optimized design sequence $\{{\boldsymbol{\xi}}_{t}^{\star}\}_{t=1:T}$ together with posterior approximations of states and parameters. BAD-PODS therefore provides a practical realization of our framework, jointly optimizing designs and performing inference online, with consistency guarantees (Section 3.4) and support for continuous design spaces. Full implementation details are given in Appendix LABEL:supp-ap:algorithm.
Algorithm 1 BAD-PODS
1:Inputs: Number of particles $(L,M,N)$, stepsizes $\{\eta_{k}\}_{k=0}^{K-1}$, prior distributions $p(\boldsymbol{\theta})$, $p(\boldsymbol{x}_{0})$.
2:Init: $\boldsymbol{\theta}_{0}^{(m)}\!\sim\!p(\boldsymbol{\theta})$, $\boldsymbol{x}_{0}^{(m,n)}\!\sim\!p(\boldsymbol{x}_{0})$; set weights uniform.
3:for $t=1,2,\dots$ do
4:  /* Design optimization (SGA) */
5:  Initialize $\boldsymbol{\xi}_{t}^{(0)}$ randomly.
6:  for $k=0$ to K1K-1 do
7:   Sample $(\widetilde{\boldsymbol{y}}_{t}^{(\ell)},\widetilde{\boldsymbol{x}}_{t}^{(m,n)},\boldsymbol{\theta}_{t-1}^{(m)})$ from $\Gamma_{\boldsymbol{\xi}_{t}^{(k)}}$.
8:   Estimate $\widehat{L}_{\boldsymbol{\theta}_{t-1},\boldsymbol{\xi}_{t}^{(k)}}^{N}$ and $\widehat{Z}_{\boldsymbol{\xi}_{t}^{(k)}}^{M,N}$.
9:   Estimate EIG gradient of (3.2).
10:   Update $\boldsymbol{\xi}_{t}^{(k+1)}\leftarrow\boldsymbol{\xi}_{t}^{(k)}+\eta_{k}\,\widehat{\nabla_{\boldsymbol{\xi}_{t}}}\mathcal{I}(\boldsymbol{\xi}_{t}^{(k)})$.
11:  end for
12:  Set $\boldsymbol{\xi}_{t}^{\star}\leftarrow\boldsymbol{\xi}_{t}^{(K)}$; collect real $\boldsymbol{y}_{t}$.
13:
14:  /* Inference: update posterior distributions*/
15:  Jitter $\boldsymbol{\theta}_{t\mid t-1}^{(m)}\sim\kappa_{M}(\cdot\mid\boldsymbol{\theta}_{t-1}^{(m)})$.
16:  Propagate $\boldsymbol{x}_{t\mid t-1}^{(m,n)}\sim f(\cdot\!\mid\!\boldsymbol{x}_{t-1}^{(m,n)},\boldsymbol{\theta}_{t\mid t-1}^{(m)},\boldsymbol{\xi}_{t}^{\star})$.
17:  Compute weights with $g(\boldsymbol{y}_{t}\!\mid\!\boldsymbol{x}_{t\mid t-1}^{(m,n)},\boldsymbol{\theta}_{t\mid t-1}^{(m)},\boldsymbol{\xi}_{t}^{\star})$.
18:  Resample first state particles, then parameter particles.
19:end for
@@ 4 RELATED WORK
A central challenge in BED is the cost and complexity of estimating the EIG (or its gradient) (Rainforth24; huan2024optimal). For models with tractable likelihoods $p(\boldsymbol{y}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi})$, the nested Monte Carlo (NMC) estimator has been studied in depth (Rainforth18), with subsequent work proposing variational formulations and bounds to improve convergence (Foster19; Foster20; Foster21). These approaches also enable differentiable objectives for gradient-based design in continuous spaces. However, they rely on explicit pointwise likelihood evaluation and do not extend directly to partially observable dynamical systems, where likelihoods require marginalization over latent states.
When likelihoods are implicit (intractable pointwise but possible to simulate from), likelihood-free design methods replace direct evaluation by surrogates or simulation-based objectives. Examples include variational mutual-information bounds and amortized estimators (Kleinegesse20; Kleinegesse21; Ivanova21), density-ratio or classifier-based estimates of information gain (Kleinegesse19), and approximate Bayesian computation (ABC)-style utilities (Dehideniya18; Drovandi13; Hainy16; Price16). While effective in simulator settings, most such methods target static design (optimize once, then deploy) rather than fully sequential/adaptive BED. Moreover, surrogate-based adaptive variants (e.g., Ivanova21) typically require an offline dataset to train an implicit likelihood, which may be unavailable or quickly misspecified in evolving dynamical systems.
Design for dynamical systems has also been studied under full state observability. Iqbal24; iqbal2024recursive consider adaptive design when the state is fully observed, making the likelihood tractable. Both papers leverage particle methods for joint state–parameter inference: (Iqbal24) uses a reversed (“inside-out”) SMC2, and (iqbal2024recursive) adopts NPF’s jittering to rejuvenate parameter particles. However, these recent advances in BED for dynamical systems assume full state observability and do not address the latent-state marginalizations required in partially observable SSMs.
@@ 5 EXPERIMENTS
We evaluate our online BAD framework on two partially observable dynamical systems. Since no existing methods target this setting, we compare against two standard BED baselines (foster2021thesis): (i) random designs sampled uniformly from $\Omega$, and (ii) static BED, a non-adaptive version of our method where the full sequence $\boldsymbol{\xi}_{1:T}$ is optimized offline. Unlike our sequential approach, the static baseline cannot adapt as data arrive.
@@
Python code to reproduce all experiments is released at https://anonymous.4open.science/r/badpods-2C8D. Details of the computing infrastructure are provided in Appendix LABEL:supp-ap:infrastructure.
@@ 5.1 Two-group SIR model
The sequential importance resampling (SIR) model partitions a population into susceptible ($S$), infected ($I$), and recovered ($R$) compartments. It has been widely used as a testbed in BED, though prior work often assumes fully observed states (Ivanova21) or static, likelihood-free design (Kleinegesse19).
We instead consider a partially observable, stochastic two-group SIR with online sequential design. Each group $g\in\{1,2\}$ has latent states $\mathbf{X}^{g}(\tau)=(S^{g}(\tau),I^{g}(\tau))^{\top}$ (the recovered $R^{g}$ being redundant since total population is fixed to $N_{g}=S^{g}+I^{g}+R^{g}$). The model parameters are infection and recovery rates $\boldsymbol{\theta}=\{(\beta_{g},\gamma_{g})\}_{g=1}^{2}$, with cross-group mixing governed by a fixed matrix $M\in\mathbb{R}^{2\times 2}$.
Dynamics follow an Itô stochastic differential equation (SDE) (allen2008introduction; ArinoVanDenDriessche2003; chatzilena2019contemporary) of the form,
$\mathrm{d}\mathbf{X}=f(\mathbf{X})\,\mathrm{d}\tau+G(\mathbf{X})\,\mathrm{d}\mathbf{W},$
with drift $f$, diffusion matrix $G$, and independent Wiener processes $\mathbf{W}$.
At each discrete time $t$, a design $\boldsymbol{\xi}_{t}=(\xi_{t}^{(1)},\xi_{t}^{(2)})$ allocates sampling effort ($\kappa>0$) across groups, constrained to the 2-simplex. Observations are noisy incident counts,
$y_{t}^{(g)}\;\big|\;\mathbf{X}_{t},\boldsymbol{\theta},\boldsymbol{\xi}_{t}\;\sim\;\mathrm{Poisson}\!\left(\kappa\,\xi_{t}^{(g)}\,\rho_{g}\,\frac{I_{t}^{g}}{N_{g}}\right),$ (9)
with known detection scales $\rho_{g}>0$. Thus $\boldsymbol{\xi}_{t}$ affects only the observation model. Simulation is performed using an Euler–Maruyama discretization. Full model details are given in Appendix LABEL:supp-ap:sir.
@@
We set the horizon to $T=200$ and simulate ground-truth trajectories from the two-group SIR model with group sizes $N_{1}\!=\!N_{2}\!=\!200$ and initial infections $I_{0}^{1}\!=\!I_{0}^{2}\!=\!5$. The cross-group mixing matrix $M$, detection scales $\rho_{g}$, sampling budget $\kappa$, and the parameters of the second group $(\beta_{2},\gamma_{2})$ are fixed and known. Inference targets only the first-group parameters $(\beta_{1},\gamma_{1})$. At each time step, the design $\boldsymbol{\xi}_{t}$ allocates observation effort between $y_{t}^{(1)}$ and $y_{t}^{(2)}$.
Our online BAD framework is implemented with a NPF using M=N=100M\!=\!N\!=\!100 particles. Design optimization at each step uses K=500K\!=\!500 stochastic gradient updates, with simplex constraints enforced via a sigmoid reparameterization. Updates are performed with Adam (DBLP:journals/corr/KingmaB14). Further details are given in Appendix LABEL:supp-ap:sir.
@@
Table 1 reports TEIG for all three methods. BAD-PODS consistently outperforms both baselines, with gaps widening over time. Results for the static method are truncated as optimization becomes infeasible at longer horizons.
Table 1: TEIG across time for two-group SIR (left) and moving source (right). Means over 50 seeds with 95% BCa CIs. Static results truncated at longer horizons due to computational cost.
Two-group SIR model Moving source model
$t$ BAD-PODS Random Static $t$ BAD-PODS Random Static
50 1.390 [1.324,1.460] 1.364 [1.287,1.445] 1.167 [1.117,1.220] 10 0.304 [0.290,0.325] 0.242 [0.222,0.263] 0.270 [0.258,0.284]
100 3.283 [3.173,3.380] 3.152 [3.040,3.268] 2.529 [2.444,2.621] 20 0.709 [0.650,0.855] 0.582 [0.531,0.649] 0.570 [0.541,0.604]
150 4.748 [4.609,4.871] 4.518 [4.385,4.627] 3.547 [3.454,3.658] 30 1.136 [1.005,1.370] 0.940 [0.855,1.044]
200 5.918 [5.765,6.055] 5.621 [5.466,5.793] 40 1.483 [1.318,1.822] 1.217 [1.136,1.344]
50 1.740 [1.551,2.096] 1.435 [1.354,1.555]
@@
Figure 1: Two-group SIR: differences in TEIG (ours minus baseline). Boxplots (median and interquartiles) compare BAD-PODS with random designs (left, green) and static BED (right, orange). Higher values indicate better performance for BAD-PODS.

### v2

@@ 3.3 Monte Carlo Estimators using NPF
Thus, to approximate the EIG in (5) we draw: (i) parameters $\{\boldsymbol{\theta}_{t}^{(m)},\dot{\boldsymbol{\theta}}_{t}^{(m)}\}$ for $m=1,\dots,M$; (ii) conditional state trajectories $\{\boldsymbol{x}_{0:t-1}^{(m,n)},\widetilde{\boldsymbol{x}}_{t}^{(m,n)},\dot{\boldsymbol{x}}_{t}^{(m,n)},\ddot{\boldsymbol{x}}_{t}^{(m,n)}\}$ for $n=1,\dots,N$; and (iii) pseudo-observations $\widetilde{\boldsymbol{y}}_{t}^{(\ell)}\!\sim\!g(\boldsymbol{y}_{t}\!\mid\!\widetilde{\boldsymbol{x}}_{t}^{(m,n)},\boldsymbol{\theta}^{(m)},\boldsymbol{\xi}_{t})$, for $\ell=1,\dots,L$. This yields the NMC approximation
@@
Following a similar procedure, the gradients $\nabla_{\boldsymbol{\xi}_{t}}L_{\boldsymbol{\theta},\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})$ and $\nabla_{\boldsymbol{\xi}_{t}}Z_{\boldsymbol{\xi}_{t}}(\boldsymbol{y}_{t})$ of (3.2) can be approximated. See details in Appendix B.
@@ 3.4 Consistency of the EIG Estimator
Consistency of our estimator $\widehat{\mathcal{I}}(\boldsymbol{\xi}_{t})$ follows from two ingredients: (i) NPF convergence under standard conditions (small/regular jittering kernels, Lipschitz state posterior in $\boldsymbol{\theta}$, and a bounded positive likelihood), and (ii) NMC consistency provided the EIG integrand, $\log(L/Z)$, is Lipschitz and square-integrable. Formal statements of these assumptions are given in Appendix C.
Theorem 3.1.
Let $\widehat{\mathcal{I}}(\boldsymbol{\xi}_{t})$ denote the NMC estimator of the EIG (9), constructed using $M$ parameter particles, NN state particles per parameter, and $L$ pseudo-observations. Under Assumptions C.2C.6 (in Appendix C), for any $t$ and $\boldsymbol{\xi}_{t}$
$\widehat{\mathcal{I}}(\boldsymbol{\xi}_{t})\;\xrightarrow[L,M,N\to\infty]{\text{a.s.}}\;\mathcal{I}(\boldsymbol{\xi}_{t}).$
The proof combines NMC results (Rainforth et al., 2018), ensuring consistency of nested expectations under Assumption C.6, with NPF convergence guarantees (Crisan and Míguez, 2018), which establish almost-sure convergence of empirical state–parameter measures under Assumptions C.2C.5. Together, these results imply that the Monte Carlo averages in our estimator converge almost surely to the true EIG. Full details are given in Appendix C.
@@ 3.5 Overall Algorithm
BAD-PODS (Bayesian Adaptive Design for Partially Observable Dynamical Systems), selects designs sequentially at each time step $t$. Given the current history $h_{t-1}$, we optimise $\boldsymbol{\xi}_{t}$ via SGA (Section 3.2), using the gradient estimator of the EIG computed with NMC samples from $\Gamma_{\boldsymbol{\xi}_{t}}$ (Section 3.3). After $K$ optimisation steps, with step sizes $\eta_{k}$, we obtain the optimized design $\boldsymbol{\xi}_{t}^{\star}$ and the real observation $\boldsymbol{y}_{t}$ is collected. The posterior is then updated via an NPF (Section 2.4): parameter particles are jittered, states are propagated through $f$, and weights are adjusted using $g$, with resampling to maintain diversity. This recursive construction avoids reprocessing past data, with a per-step cost of $\mathcal{O}((KL+1)MN)$ (and $\mathcal{O}(KLMNt)$ over $t$ steps), ensuring linear scaling in time while supporting gradient-based design optimization.
The algorithm takes as input the prior distributions, particle numbers $(M,N)$, and optimization hyperparameters $(K,\eta_{k})$, and produces as output the optimized design sequence $\{{\boldsymbol{\xi}}_{t}^{\star}\}_{t=1:T}$ together with posterior approximations of states and parameters. See Appendix D for full implementation details.
@@ 4 Related Work
A central challenge in BED is the cost and complexity of estimating the EIG (or its gradient) (Rainforth et al., 2024; Huan et al., 2024). For models with tractable likelihoods $p(\boldsymbol{y}\!\mid\!\boldsymbol{\theta},\boldsymbol{\xi})$, the NMC estimator has been studied in depth (Rainforth et al., 2018), with subsequent work proposing variational formulations and bounds to improve convergence (Foster et al., 2019, 2020, 2021). These approaches also enable differentiable objectives for gradient-based design in continuous spaces. However, they rely on explicit pointwise likelihood evaluation and do not extend directly to partially observable dynamical systems, where likelihoods require marginalization over latent states.
When likelihoods are implicit (intractable pointwise but possible to simulate from), likelihood-free design methods replace direct evaluation by surrogates or simulation-based objectives. Examples include variational mutual-information bounds and amortized estimators (Kleinegesse and Gutmann, 2020, 2021; Ivanova et al., 2021), density-ratio or classifier-based estimates of information gain (Kleinegesse and Gutmann, 2019), and approximate Bayesian computation (ABC)-style utilities (Dehideniya et al., 2018; Drovandi and Pettitt, 2013; Hainy et al., 2016; Price et al., 2016). While effective in simulator settings, most such methods target static design (optimize once, then deploy) rather than fully sequential/adaptive BED. Moreover, surrogate-based adaptive variants (e.g., Ivanova et al., 2021) typically require an offline dataset to train an implicit likelihood, which may be unavailable or quickly misspecified in evolving dynamical systems.
Design for dynamical systems has also been studied under full state observability. Iqbal et al. (2024b, a) consider adaptive design when the state is fully observed, making the likelihood tractable. Both papers leverage particle methods for joint state–parameter inference: Iqbal et al. (2024b) use a reversed (“inside-out”) SMC2, and Iqbal et al. (2024a) adopt NPF’s jittering to rejuvenate parameter particles. However, these recent advances in BED for dynamical systems assume full state observability and do not address the latent-state marginalizations required in partially observable SSMs.
@@ 5 Experiments
We evaluate our method on two partially observable dynamical systems. Since no existing methods target this setting, we compare against three standard BED baselines (Foster, 2021): (i) random designs sampled uniformly from $\Omega$, (ii) an oracle version of BAD-PODS, which (instead of gradient-based optimisation) evaluates the estimated EIG on a dense discretisation of $\Omega$ and selects the best grid point at each step (providing an approximate upper bound); and (iii) static BED, a non-adaptive version of our method where the full sequence $\boldsymbol{\xi}_{1:T}$ is optimized offline. Unlike our sequential approach, the static baseline cannot adapt as data arrive.
@@
Python code to reproduce all experiments is released at https://anonymous.4open.science/r/badpods-2C8D. Details of the computing infrastructure are provided in Appendix H.
@@ 5.1 Two-group SIR Model
The susceptible-infectious-recovered (SIR) model is a standard testbed in BED, but prior work often assumes fully observed states (Ivanova et al., 2021) or static design (Kleinegesse and Gutmann, 2019). Here we use a partially observable, stochastic two-group SIR model with online sequential design. The latent state tracks the susceptible/infectious counts in each group $\mathbf{X}^{(g)}(\tau)=\big(S^{(g)}(\tau),I^{(g)}(\tau)\big)^{\top}$, and the unknown parameters are the infection and recovery rates $\boldsymbol{\theta}=\{(\beta^{(g)},\gamma^{(g)})\}_{g=1}^{2}$. At each time step $t$, the design $\boldsymbol{\xi}_{t}=(\xi_{t}^{(1)},\xi_{t}^{(2)})$ allocates a fixed sampling effort across the two groups (simplex constraint), producing noisy incidence observations. The design therefore modulates the observation process, while the epidemic dynamics remain unchanged. Full model details are given in Appendix E.
@@
We set the horizon to T=200T\!=\!200 and simulate ground-truth trajectories from the two-group SIR model with group sizes $N^{(1)}\!=\!N^{(2)}\!=\!200$ and initial infections $I_{0}^{(1)}\!=\!I_{0}^{(2)}\!=\!5$. The cross-group mixing matrix $M$, detection scales $\rho^{(g)}$, sampling budget $\kappa$, and the parameters of the second group $(\beta^{(2)},\gamma^{(2)})$ are fixed and known. Inference targets only the first-group parameters $(\beta^{(1)},\gamma^{(1)})$. At each time step, the design $\boldsymbol{\xi}_{t}$ allocates observation effort between $y_{t}^{(1)}$ and $y_{t}^{(2)}$. For more details see Appendix E.
@@
Table 1 reports TEIG for all four methods. BAD-PODS consistently performs close to the oracle solution, and outperforms both static and random baselines, with gaps widening over time. Results for the static method are truncated as optimization becomes infeasible at longer horizons.
Table 1: TEIG across time for the two-group SIR model (top), moving-source model (middle) and ecological growth model (bottom). Means over 50 seeds with 95% BCa CIs. Static results truncated at longer horizons due to computational cost.
Two-group SIR model
$t$ Oracle BAD-PODS Random Static
50 1.450 1.390 1.364 1.167
[1.376,1.520] [1.324,1.460] [1.287,1.445] [1.117,1.220]
100 3.330 3.283 3.152 2.529
[3.215,3.442] [3.173,3.380] [3.040,3.268] [2.444,2.621]
150 4.848 4.748 4.518 3.547
[4.747,4.959] [4.609,4.871] [4.385,4.627] [3.454,3.658]
200 6.043 5.918 5.621
[5.939,6.168] [5.765,6.055] [5.466,5.793]
Moving source model
$t$ Oracle BAD-PODS Random Static
10 0.304 0.297 0.242 0.270
[0.296,0.312] [0.289,0.308] [0.222,0.263] [0.258,0.284]
20 0.698 0.688 0.582 0.570
[0.672,0.731] [0.661,0.726] [0.531,0.649] [0.541,0.604]
30 1.158 1.130 0.940
[1.103,1.223] [1.075,1.197] [0.855,1.044]
40 1.515 1.465 1.217
[1.451,1.591] [1.400,1.556] [1.136,1.344]
50 1.805 1.733 1.435
[1.732,1.888] [1.658,1.840] [1.354,1.555]
Ecological growth model
$t$ Oracle BAD-PODS Random Static
5 0.935 0.902 0.725 0.897
[0.896, 0.970] [0.854, 0.946] [0.657, 0.790] [0.866, 0.946]
10 1.618 1.579 1.337 1.571
[1.554, 1.690] [1.499, 1.650] [1.245, 1.429] [1.507, 1.627]
15 2.277 2.215 1.972
[2.175, 2.379] [2.113, 2.312] [1.844, 2.109]
20 2.621 2.545 2.344
[2.492, 2.741] [2.422, 2.673] [2.211, 2.499]
@@
[…]
@@
Runtime and memory.
For the SIR experiment with T=150T\!=\!150, average wall-clock runtimes are $2.09$ minutes for the random baseline (no optimisation), $31.05$ for the static baseline (single high-dimensional optimisation over $\boldsymbol{\xi}_{1:T}$), and $12.39$ for BAD-PODS (per-step optimisation with online inference), yielding a $\approx\!2.5\times$ speed-up over the static approach. Memory is an equally important bottleneck: at T=150T\!=\!150, the static baseline uses $\approx 25$ GB GPU and $\approx 25$ GB CPU memory, whereas BAD-PODS uses $\approx 10$ GB GPU and $\approx 1$ GB CPU. This gap grows with $T$ since BAD-PODS only needs to store the current $M\!\times\!N$ particle set, while the static baseline must store and optimise over the full design tensor and latent trajectories of size $M\!\times\!N\!\times\!T$, making long horizons increasingly expensive.